If I have a diagonal of 220 and a width of 150, what is the height of the rectangle?
Answer by unlockmath(1688)

Hello,
We can set up this using this formula.
A^2 + B^2 = C^2
C is the diagonal and width is A.
150^2 + B^2 = 220^2
Expand this out to:
22500 + B^2 = 48400
Subtract 22500 and we get:
B^2 = 25900
Square root both sides.
B = +-160.935
Only +160.935 works.
Make sense?
RJ
www.math-unlock.com